• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 16:06:14 on August21,2003

    Entry number 112437

    keyword=Shift Summary (day shift)


    From 08:00 - 08:45
    Took limited statistics in kinematics 10.20 due to pass change.

    At MCC's request and after checking with RC we turned off
    Q1 on both spectrometer and set the septum to 0 current for
    RadCon tests.

    Did beam current calibration followed by septum heating test.

    Max. beam current from the injector was limited to 80 uA.

    We saw approx. 0.5k heating of the upper and lower septum coils
    with the multifoil target in an 80 uA of beam.